OpenStudy (anonymous):

how do you calculate the following determinant: det((-2B^T)^-3)) when det(B)=5

OpenStudy (amistre64):

since the determinant doesnt care if its Ted or not; B^T = 5 the ^-3 might have me concerned tho

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/LinAlg/DeterminantProperties.aspx this might be useful to read up on

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

its still that ^-3 thats got me a little concerned; you sure you typed it correctly?

OpenStudy (anonymous):

yeah that means that its 1/(-2B^t)^3 but i dunno where to go from there. like if there's a property to help or smthg

OpenStudy (amistre64):

det(AB) = det(A) det(B) if A and B are the same size

OpenStudy (anonymous):

okayy but what do you do when the determinant is to the 3rd power?

OpenStudy (amistre64):

wht does the third power mean?

OpenStudy (anonymous):

like 2nd power is squared, third power....

OpenStudy (anonymous):

cubed

OpenStudy (amistre64):

well, we know the thrm for det(AB) is just multiplications ... are powers multiplications?

OpenStudy (anonymous):

well its a multiplication by itself. so the determinant multiplied by itself 3 times

OpenStudy (amistre64):

id say so yes. that should be all the thrms we need to solve it them

OpenStudy (anonymous):

aightosss amigosssssss danke freund :)

OpenStudy (amistre64):

bitteshun .. if i recall me duetch \[\frac{1}{(-2*5)^3}\]

OpenStudy (anonymous):

right so what do u do after that? is it like 1/ (-2)^3*5^3

OpenStudy (anonymous):

btw ur deutsch is fantabolical

OpenStudy (zarkon):

what is the size of B

OpenStudy (anonymous):

its a n3X3

OpenStudy (amistre64):

ja ...meine grossmutter sint duetch

OpenStudy (zarkon):

that matters

OpenStudy (anonymous):

how so?

OpenStudy (zarkon):

\[Det(cA)=c^nDet(A)\] when \(A\) is an \(n\times n\) matrix

OpenStudy (amistre64):

thrm 1 ....yeah

OpenStudy (zarkon):

missed your link :)

OpenStudy (anonymous):

omgooshhh thanks :) i could kiss you right now